Circular No. 8494

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 E2 (McNAUGHT) R. H. McNaught reports his discovery of a comet with a 12′′ tail in

p.a. 250^{o} on CCD images taken on Mar. 12.8 and 13.8 UT with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation given below). Following posting on the ‘NEO Confirmation Page’, CCD observations obtained with the 1.0-m f/7.7 reflector at Mt. John University Observatory on Mar. 13.7 by A. C. Gilmore (measured by P. M. Kilmartin) show a uniform coma ∼ 7′′ in diameter with a stubby tail 13′′ long to the southwest.

2005 UT  α_{2000}  δ_{2000}  Mag.
Mar. 12.75009  19^{h}47^{m}24.^{s}43  −35^{o}48′32′′.6  16.5

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-E84.

T = 2006 Apr. 13.126 TT  ω = 358.^{o}748
Ω = 357.427 }2000.0
q = 3.01128 AU  i = 16.116

COMETS C/2005 E3 (SOHO) AND C/2005 E4 (SOHO) Additional comets found on SOHO website images (cf. IAUC 8490;

MO = M. Oates).

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 E3 Mar. 6.671 23^{h}16.^{m}9 − 6^{o}09′ C2 MO 2005-E87
C/2005 E4  10.388 23 33.0 − 2 48 C3 RK 2005-E87

C/2005 E3 is a Kreutz sungrazer, while C/2005 E4 belongs to the Marsden group. On MPEC 2005-E87, B. G. Marsden suggests that C/2005 E4 is identical with C/1999 N5 and has P = 5.66 yr. He shows that, when run back a further revolution, the linked orbit agrees very closely with the linked orbit corresponding to his earlier suggestion that C/2004 V9 was a return of C/1999 J6 (IAUC 8456, MPEC 2004-X73, P = 5.49 yr), with then T = 1993 Nov. 20–22, q = 0.0515 AU, and agreement to better than 0^{o}.3 in ω, Ω, and i. This indicates that these Marsden-group comets probably separated from each other around that time. Marsden adds that, if C/1999 U2 also then separated from them, it would now have P = 5.95 yr and T = 2005 Oct. 8 (to within a few days).

2005 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green